Fermi National Laboratory
 

Operation Summary for December 3, 2010 - December 6, 2010

 

The Integrated Luminosity for the period from 11/29/10 to 12/6/10 was 61.97 pb-1 (inverse picobarns). NuMI reported receiving 8.1E18 protons on target during this same period.

Friday Dec 3

The day shift began with Operations monitoring the Tevatron (TeV) store 8323 (83.37E30) and the Recycler stash (331.2E10), with the Antiproton Source (Pbar) stacking (2.4E10), with Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson FTBF T-1008 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations reported that a Linac RF station (LRF1) was down due to a power amplifier vacuum problem. Linac technicians had it repaired by 9:43 AM. Operations terminated store 8323 at 10:06 AM. Kicker personnel completed their work on the TeV abort kickers and TeV personnel began a study period. After the studies finished, operators prepared TeV sectors A-E and the CDF and D0 collision halls for access. Cryo system techs accessed the TeV to work at D1 and for the C3 wet engine repair. Vacuum techs accessed to work through out all the TeV sectors. EE Support personnel repaired a P2 line correction magnet bulk power supply; this created about 45 minutes of stacking downtime. All three LCW pumps at MI-10 tripped off. FESS personnel investigated.

During the evening shift, Operations reported that all the TeV repairs were completed by 6:30 PM. However, during the TeV checkout, operators found a low beta quadrupole magnet that had a ground fault indication; operators called in EE Support personnel to investigate. EE Support personnel reported finding LCW leaks on three of the four low beta power supplies; they had all the leaks repaired by 11:45 PM. Operations began the TeV turn on. MI personnel decided to run without the MI LCW pumps until all of the antiprotons in the Recycler were transferred to the TeV during a shot so as to not risk losing them when the pumps were turned on.

Saturday Dec 4

The midnight shift began with the TeV turn on in progress, with Operations monitoring the stash (71.7E10), with Pbar stacking (7.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators completed the TeV turn on at 4:08 AM and Operations put the TeV into HEP shot setup. Operations established store 8330 at 4:59 AM with an average initial luminosity (AIL) of 302.67E30. Operations established beam to T-1008 at 5:52 AM; Operations reported problems in getting the desired intensity to the experiment and were tuning.

During the day shift, operators continued to tune beam to the T-1008 experiment. Operations terminated store 8330 at 1:04 PM, put the TeV into shot setup, and established store 8331 at 2:13 PM with an AIL of 323.6E30. After the shot, operators turned on the MI-10 LCW pumps, which resulted in a half hour of downtime.

During the evening shift, Operations reported that the MI-52 LCW pump tripped off, which caused another hour of downtime.

Sunday Dec 5

The midnight shift began with Operations monitoring store 8331 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 1:04 AM, store 8331 aborted due to vacuum problems at C4. Operators replaced a bad crate power supply, however the crate would still not produce any readbacks. Operators called in Controls personnel and they found a bad connector on the back plane. Operations established beam to T-1008 at 4:33 AM. Operators recovered the TeV and put it into shot setup at 6:44 AM. Operations established store 8332 at 7:39 AM with an AIL of 333.37E30. Operations reported multiple trips of LRF2; operators raised the driver forward power to stabilize the station.

During the day shift, operators continued to tune beam to the T-1008 experiment.

During the evening shift, Operations terminated store 8332 at 6:03 PM. Operators prepared CDF for an access that ended at 6:59 PM. Operators began the TeV turn on and orbit checks. Operators and Switchyard (SY) personnel continued to tune the beam to T-1008. Operations reported a short downtime due to the Booster RF systems. Operations put the TeV into shot setup at 8:30 PM and established store 8335 at 9:39 PM with an AIL of 231E30. The luminosity suffered from a separator spark during the low beta squeeze.

Monday Dec 6

The midnight shift began with Operations monitoring store 8335 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 1:01 AM, store 8335 quenched due to a broken crankshaft for the C1 wet engine; the wet engine also needed a flywheel replacement. Cryo system techs completed the work and operators had recovered the TeV by 6:12 AM. Operators began the TeV checkout. Operations put the TeV into shot setup at 7:32 AM. Operations reported that the BRF12 modulator tripped off and required several minutes to reset.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Electron Cooling personnel to conduct a study and for MI personnel to try and improve the beam to SY and the T-1008 experiment.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/06/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory